If C(n,3)=C(n,4),then the value of n?????please help me to solve this problem.
Answer by macston(5194)

.
C(n,3)=n!/(3!(n-3)!)
.
C(n,4)=n!/(4!(n-4)!)
.
n!/(3!(n-3)!)=n!/(4!(n-4)!) Divide each side by n!
1/(3!(n-3)!)=1/(4!(n-4)!) Cross multiply
3!(n-3)!=4!(n-4)! Divide each side by 3!
(n-3)!=4(n-4)! Divide each side by (n-4)!*(Note below)
n-3=4(1)
n-3=4
n=7
ANSWER: n=7
.
*NOTE: (n-3)!=(n-3)(n-4)! so (n-3)!/(n-4)!=(n-3)
.
CHECK:
n!/(3!(n-3)!)=n!/(4!(n-4)!)=
7!/(3!(7-3!))=7!/(4!(7-4)!)=
7!/(3!4!)=7!/(4!3!)
(7*6*5)/(3*2*1)=(7*6*5)/(3*2*1)
210/6=210/6
35=35